Forrester Company is considering buying new equipment that would increase monthly fixed costs from $130,000 to $150,000 and would decrease the current variable costs of $70 by $10 per unit. The selling price of $120 is not expected to change. Forrester's current break-even sales are $312,000 and current break-even units are 2,600. If Forrester purchases this new equipment, the revised break-even point in units would:

Multiple Choice Increase by 100. Decrease by 100. Increase by 13,000. Decrease by 10,400. Increase by 10,400.
Answer #1

New fixed cost = 150,000

variable cost = 10

Selling price = 120

150,000 + (60 x units) = 120 x units

150,000 = (120-60) x units

150,000 = 60 units

2500 units

Previous break even point (units) = 2600

Difference = 2600 - 2500

= 100

Decrease by 100